# Parity violation as enforced symmetry breaking in 3D fermionic topological order

### v1

@@ Appendix D Topological crystalline phases corresponding to the compatible case
Here we discuss how to construct the topological crystalline phases that corresponding to the $\mathbb{Z}_{4}^{\mathcal{T}f}\rtimes\mathbb{Z}_{8}\times\mathbb{Z}_{2}$ fSPT that can reduce to the $\mathbb{Z}_{2}^{f}\times\mathbb{Z}_{8}\times\mathbb{Z}_{2}$ fSPT featuring non-Abelian three loop statistics. We choose to transform to $\mathbb{Z}_{2}$ subsymmetry into the $C_{2}$ rotation symmetry, so that the topological crystalline phases are protected by $\mathbb{Z}_{4}^{\mathcal{T}f}\rtimes\mathbb{Z}_{8}\times C^{-}_{2}$ symmetry, where $C_{2}^{-}$ here we mean the square of the $C_{2}$ rotation equals to fermionic parity [41]. To see the construction, we first make two remarks. The first one is that the cell decomposition for $C_{2}$ rotation system in 3D consists of a 1D spatial line as the rotation center and two 2D planes that intersect at the rotation center and can be transformed into each other by rotation, and two 3D cells that are also transformed into each other by rotation. To construct $\mathbb{Z}_{4}^{\mathcal{T}f}\rtimes\mathbb{Z}_{8}\times C^{-}_{2}$ symmetric topological phases, we can place 1D, 2D, or 3D topological phases with certain symmetry on the 1D, 2D, or 3D cell accordingly. The second remark is that the classification of 2D $\mathbb{Z}_{4}^{\mathcal{T}f}\rtimes\mathbb{Z}_{8}$ fSPT is summarized in Table.4 where we can see the classification from KC decoration is $(\mathbb{Z}_{2})^{2}$: one $\mathbb{Z}_{2}$ is purely protected by TRS with $\mathcal{T}^{2}=\mathcal{P}_{f}$ while the other is purely by $\mathbb{Z}_{8}$. The decoration of 2D $\mathbb{Z}_{4}^{\mathcal{T}f}\rtimes\mathbb{Z}_{8}$ fSPT characterized by the purely $\mathbb{Z}_{8}$ KC decoration on the 2D cell gives the topological crystalline phases we are looking for. As such a 2D state is $\mathbb{Z}_{2}$ classified, so along the rotation line there are two copies of 1D edge theory of such a 2D fSPT, which together is ’t Hooft anomaly-free, so that they can be symmetrically gapped out to complete the crystalline phase construction.

### v2

@@ Appendix A Topological crystalline phases corresponding to the compatible case
Here we discuss how to construct the topological crystalline phases that corresponding to the $\mathbb{Z}_{4}^{\mathcal{T}f}\rtimes\mathbb{Z}_{8}\times\mathbb{Z}_{2}$ fSPT that can reduce to the $\mathbb{Z}_{2}^{f}\times\mathbb{Z}_{8}\times\mathbb{Z}_{2}$ non-Abelian fSPT. We choose to transform to $\mathbb{Z}_{2}$ subsymmetry into the $C_{2}$ rotation symmetry, so that the topological crystalline phases are protected by $\mathbb{Z}_{4}^{\mathcal{T}f}\rtimes\mathbb{Z}_{8}\times C^{-}_{2}$ symmetry, where $C_{2}^{-}$ here we mean the square of the $C_{2}$ rotation equals to fermionic parity. To see the construction, we first make two remarks. The first one is that the cell decomposition for $C_{2}$ rotation system in 3D consists of a 1D spatial line as the rotation center and two 2D planes that intersect at the rotation center and can be transformed into each other by rotation, and two 3D cells that are also transformed into each other by rotation. To construct $\mathbb{Z}_{4}^{\mathcal{T}f}\rtimes\mathbb{Z}_{8}\times C^{-}_{2}$ symmetric topological phases, we can place 1D, 2D, or 3D topological phases with certain symmetry on the 1D, 2D, or 3D cell accordingly. The second remark is that the classification of 2D $\mathbb{Z}_{4}^{\mathcal{T}f}\rtimes\mathbb{Z}_{8}$ fSPT is summarized in Table.4 where we can see the classification from KC decoration is $(\mathbb{Z}_{2})^{2}$: one $\mathbb{Z}_{2}$ is purely protected by TRS with $\mathcal{T}^{2}=\mathcal{P}_{f}$ while the other is purely by $\mathbb{Z}_{8}$. Now we claim that the decoration of 2D $\mathbb{Z}_{4}^{\mathcal{T}f}\rtimes\mathbb{Z}_{8}$ fSPT characterized by the purely $\mathbb{Z}_{8}$ KC decoration on the 2D cell gives the topological crystalline phases we are looking for. As such a 2D state is $\mathbb{Z}_{2}$ classified, so along the rotation line there are two copies of 1D edge theory of such a 2D fSPT, which together is ’t Hooft anomaly-free, so that they can be symmetrically gapped out to complete the crystalline phase construction.
We also note that another compatible example exists with the TRS action $\mathbf{r}=(1,0,1)$. We further note that two scenarios with automorphisms $\mathbf{r}=(3,0,0)$ and $(3,0,1)$ are consistent with this non-Abelian TO and are both isomorphic to $G_{f}=(\mathbb{Z}_{4}^{Tf}\times\mathbb{Z}_{2}\times D_{16})/\mathbb{Z}_{2}^{T}$.